OpenStudy (anonymous):

Hello! I need help with the following question please ( I will post as a comment so that I can use the equation thing)

OpenStudy (anonymous):

\[(\sqrt{2x}-y)(\sqrt{2x}+y)-(4x-y)^{2}\]

OpenStudy (anonymous):

The question asks me to simplify

OpenStudy (anonymous):

But I dont get the right answer when I try it. I get \[-14x ^{2}+8xy\]

hartnn (hartnn):

show your work ?

OpenStudy (anonymous):

So when I expand the first one I get\[2x ^{2} + \sqrt{2xy} -\sqrt{2xy}+y ^{2} -(4x -y)^{2}\]

OpenStudy (anonymous):

Which I then simplify to\[2x ^{2} +y ^{2} - 16x ^{2}-4xy-4xy-y ^{2}\]

hartnn (hartnn):

the first term shouldn't be 2x^2

hartnn (hartnn):

\(\sqrt a \times \sqrt a = a\) \(\sqrt {2x} \times \sqrt {2x} = ??\)

OpenStudy (anonymous):

Oh! Should it be 4x^2 ?

hartnn (hartnn):

its square root

OpenStudy (anonymous):

No no thats not right..

hartnn (hartnn):

\( 2x \times 2x = 4x^2 \)

OpenStudy (anonymous):

Ok

hartnn (hartnn):

\(\sqrt {2x} \times \sqrt {2x} = ??\)

hartnn (hartnn):

\(\sqrt ☻ \times \sqrt ☻ = ☻\)

OpenStudy (anonymous):

I still get 2x^2

hartnn (hartnn):

how did you get the exponent of x as 2 ?

OpenStudy (anonymous):

oooh. I thought since x multiplied by x.. but its part of the square root right? so the exponent would still be one?

hartnn (hartnn):

\(\sqrt a \times \sqrt a = \sqrt {a^2 } = a\)

hartnn (hartnn):

yes

OpenStudy (anonymous):

Hooray! Was that my only mistake?

hartnn (hartnn):

\( (+y) \times (-y) = ... ?\)

OpenStudy (anonymous):

= -y^2

hartnn (hartnn):

i see you wrote it as +y^2

OpenStudy (anonymous):

Oh, so they are both negative? the two y^2 ?

hartnn (hartnn):

no, \(-(4x-y)^2\) lets work on this separately

OpenStudy (anonymous):

Ok

hartnn (hartnn):

your attempt was \(-16x^2 −4xy−4xy−y^2\) did you distribute the negative sign to ALL the expanded terms of \((4x-y)^2\) ?

hartnn (hartnn):

oh and yes, both the y^2 's are negatives.

OpenStudy (anonymous):

Ok, No I didnt distribute, I thought it could be left alone which I realize is not at all correct

hartnn (hartnn):

yes. to avoid confusion, i sugges you retain the brackets while expanding...anything

OpenStudy (anonymous):

I see, thank you

hartnn (hartnn):

\(-(a-b)^2 = -(a^2 -2ab+b^2 ) = -a^2 +2ab - b^2 \)

OpenStudy (anonymous):

So let me try it with the distribution..

hartnn (hartnn):

sure, let me know your final answer, i think this time you'll get it :)

OpenStudy (anonymous):

\[-16x ^{2} + 8xy -y ^{2}\]

hartnn (hartnn):

there were 2 y^2's which were negative, right? -y^2 - y^2 = ...? also add the first term, 2x in that

hartnn (hartnn):

\(\sqrt {2x} \times \sqrt {2x} = 2x\)

OpenStudy (anonymous):

Yes, I got positive y^2 and then outside negative sign made it negative

hartnn (hartnn):

ok good, so we now have \(2x -y^2 -16x^2 +8xy -y^2\) right? just one simplification, and we're done

OpenStudy (anonymous):

-16x^2 + 8xy +2x -2y^2

hartnn (hartnn):

correct!

OpenStudy (anonymous):

Thank you very much!

hartnn (hartnn):

welcome ^_^